OpenStudy (newwar):

To get on the top players’ list, Don needs to have a minimum average score of 225 after playing four games. His scores on his first three games were 192, 214, and 250. What is the minimum score Don needs to earn on his fourth game? Enter your answer in the box.

OpenStudy (newwar):

@chpatterson @3mar @Candystarlover2

OpenStudy (3mar):

One minute please

OpenStudy (newwar):

OK THX

OpenStudy (newwar):

@3mar

OpenStudy (3mar):

Sorry for late!

OpenStudy (newwar):

its ok

OpenStudy (3mar):

Sorry for late. It was an emergency call.

OpenStudy (3mar):

This problem is related to mean value or arithmetic mean.

OpenStudy (3mar):

It is calculated with the form: For a, b, c, d, e, and f, the arithmetic mean is\[\frac{ a+b+c+d+e+f }{ 6 }\] where 6 represents the number of items you are calculating the arithmetic mean for.

OpenStudy (3mar):

For our case: calculate the arithmetic mean for the first three games; we get \[\frac{ 192+214+250 }{ 3 }=218.667\] then he needs to win the fourth game with minimum score x to get the average score of 225 so the arithmetic mean of x and 218.66 will be 225 \[\frac{ 218.667+x }{ 2 }=225\] \[x=231.333\] so the fourth game score should be at least 232